# SplatCo: Structure-View Collaborative Gaussian Splatting for Detail-Preserving Rendering of Large-Scale Unbounded Scenes

### v1

@@ Abstract
Index Terms:
3DGS, real-time rendering, large-scale unbounded scenes, detail preservation, structure-view collaborative learning.
@@ I Introduction
In this work, we propose Structure–View Collaborative Gaussian Splatting (SplatCo), a unified framework for high-fidelity reconstruction and rendering of large-scale outdoor scenes. Unlike prior approaches that separately address geometry representation or view consistency, SplatCo integrates both global structure modeling and multi-view coherence into a single, jointly optimized system. First, we incorporate multi-scale contextual information into a global tri-plane backbone using a three-level compensation mechanism, which aligns coarse scene layout with detailed local geometry and texture. Second, to maintain consistency across viewpoints, we synchronize gradient updates across overlapping views, ensuring coherent Gaussian positioning and appearance. Third, we introduce a visibility-aware optimization process that densifies sparsely observed regions while removing overfitted or inconsistent Gaussians based on structural consistency. By combining these components in a unified training and inference pipeline, SplatCo achieves robust generalization and fine-detail reconstruction across complex, large-scale environments, as shown in Fig. 1.
@@
  • We introduce a cross-view assisted training strategy that synchronizes gradient updates across overlapping views, reducing single‑view overfitting and improving multi-view consistency in large-scale scenes.
@@ II-B Novel View Synthesis
3DGS has been widely adopted for modeling objects and scenes due to its speed and rendering quality. Readers may refer to recent surveys for comprehensive overviews of 3DGS and its applications [47, 48, 49]. While early applications focused on compact scenes, such as individual objects [50, 51], humans [52, 53, 54], and indoor environments [55, 56, 57], adapting 3DGS to large-scale outdoor scene remains a challenge.
To bridge this, recent works propose scalable variants. Scaffold-GS [58] introduces a hierarchical, region-aware 3D representation via anchor points. DOGS [59] significantly improves training efficiency through a distributed paradigm that combines spatial partitioning with the ADMM solver. VastGaussian [60] and Hierarchy-GS [61] explore alternative partitioning strategies, each offering distinct architectural advantages. More recent developments include Octree-GS [12], which organizes Gaussian primitives in an octree structure to support adaptive multi-scale rendering, and CityGaussian [62], which incorporates level-of-detail techniques to optimize both training and rendering in urban-scale environments.
Despite these advances, two core limitations remain. First, rigid partitioning of space often breaks continuity of semantic structures, such as roads, facades or vegetation. Second, the lack of structural coordination and view-collaboration mechanisms limits the rendering fidelity required in large-scale scenarios. These challenges often , lead to visual artifacts, such as ghosting, texture flickering, and inconsistent shading.
Figure 2: Overview. We propose a Cross-Structure Collaboration Module (CSCM) that hierarchically integrates tri-planes with context grids, effectively preserving fine-grained details while maintaining consistent global geometry. Additionally, we introduce a Cross-View Assisted Training Strategy (CVATS) that enforces gradient consistency across overlapping cameras, effectively mitigating single-view overfitting in large-scale scenes. To further enhance generalization, we develop a structural-consistency-guided pruning mechanism, which iteratively removes over-specialized Gaussians, thereby preserving geometric integrity and ensuring robust performance across diverse viewpoints.
@@ III-C Cross-Structure Collaboration Module
Tri-plane Feature Acquisition. Tri-Plane is a static 3D instance of the $k$-planes factorization [64], which represents the 3D space using three $m$-channel feature plane maps $\mathbf{F}\in\mathbb{R}^{w\times h\times m}$, each corresponding to one of the axis-aligned planes: $xy$, $xz$, and $yz$, with spatial resolution $w\times h$. To obtain fine-grained tri-plane features for a queried 3D point $\mathbf{p}$, we first apply a tri-plane attention module to generate attention-enhanced tri-plane features $\mathbf{F}^{TA}$. This module captures relative features across all three planes from the global tri-plane context, thereby enhancing the global perception capability compared to traditional tri-plane representations that focus solely on individual planes. The formulation is as follows:
@@
where $i$ denotes the axis-aligned planes: $xy$, $xz$, and $yz$, and $\text{TA}(\cdot)$ represents the tri-plane attention operation. The operator $||_{j\in\{xy,xz,yz\}}(\cdot)$ indicates concatenating feature planes along the channel dimension in the order of $xy$, $xz$, and $yz$. $\text{SA}(\cdot)$ and $\text{CA}(\cdot)$ denote the spatial attention and channel attention operations [65] applied to the feature maps, respectively. $\rho_{i}(\cdot)$ corresponds to the operation of extracting the tri-plane attention feature map for the dimension indexed by $i$ from the previously concatenated feature maps.
@@
[…]
@@ III-D Cross-View Assisted Training Strategy
To address this issue, we propose a Cross-View Assisted Training Strategy (CVATS). Specifically, we first introduce a cross-view stereo regularization method that leverages multiple views for collaborative optimization. Additionally, we propose cross-view augmented densification to improve point distribution by incorporating information from spatially distant views, which is particularly advantageous in large-scale scenes. Finally, a cross-view consistent loss strategy is employed to enforce coherence between adjacent viewpoints and mitigate overfitting of Gaussians.
Cross-View Stereo Regularization. Unlike the original single-view iterative training, we introduce a cross-view stereo approach that optimizes 3D Gaussians with multi-view supervision. In each iteration, we sample $M$ pairs of images and camera parameters, ensuring they are distinct from one another. Our multi-view regulated learning, which integrates gradients in a single iteration, is formulated as:
@@
where $\mathcal{L}$ represents the loss at the current iteration. Let $G=\{G_{1},G_{2},\dots,G_{M}\}$ denote the set of 3D Gaussians, with each $G_{M}$ being affected by large gradients from each view during multi-view training. The extrinsic and intrinsic parameters of the camera are denoted by $E_{M}$ and $K_{M}$, respectively.
@@
Cross-View Augmented Densification. When optimizing across multiple viewpoints, particularly in large-scale scenes, selecting views that are too far apart is common. This results in limited overlap between perspectives, which increases the risk of overfitting Gaussian ellipsoids. To address this, we first normalize the camera translations of sampled views to fit within a unit sphere, making our method adaptable to various scenes. We then compute the relative translation distances $\{r_{i}\mid i=1,2,\dots,M\}$ between the cameras. Our multi-view augmented densification employs a self-adaptive threshold $\hat{\beta}$, which is formulated as:
@@
Mega-NeRF [CVPR22] [66] - - - 0.553 24.06 0.516 0.547 20.93 0.504
Switch-NeRF [ICLR22] [67] - - - 0.562 24.31 0.496 0.579 21.54 0.474
@@
III-E Update Gaussian Attributes &\&& Rasterization
@@
where $\mathcal{A}_{i}$ represents the parameters of the $k$ Gaussian primitives for each anchor, and []delimited-[][$\cdot$][ ⋅ ] denotes the concatenation operation.
@@ IV-C Fine-Grained Gaussian Regularization Loss
where $N_{g}$ represents the number of neural Gaussians in the scene, and ($\prod$($\cdot$)∏ ( ⋅ ) indicates the product of the elements in a vector. In this case, it applies to the scale values $s_{i}$ of each Gaussian.
@@ V-A Datasets
Self-collected Scenes. To further evaluate the performance of our method in campus environments and plateau regions with large coverage areas and rich details, we collect the SCUT Campus Aerial (SCUT-CA) scenes, which consist of three scenes featuring teaching, office, and history museum buildings. Specifically, we use the DJI Matrice 300 RTK to capture images from approximately 100 meters above three distinct scenes at the Wushan Campus of South China University of Technology, with 125, 197, and 273 images, respectively, totaling 595 images. Each image has a resolution of 4082$\times$2148. To maintain consistency with the 3DGS [7] input, we downsample the images to a resolution of 1600$\times$841.
We also collect scenes from plateau regions, which include two scenes characterized by vast distribution, along with diverse geographical features. Specifically, we use the DJI Matrice 300 to capture images at an altitude of approximately 100 meters in the plateau region (Linzhi, Tibet), obtaining 203 and 170 images, respectively, for a total of 373 images. Each image has a resolution of 6222$\times$4148, and we downsample these images to a resolution of 1600$\times$841 as well. Camera parameters are estimated using widely adopted techniques such as COLMAP [76].
@@ V-D1 Result on Mill-19 and MatrixCity dataset
We rigorously evaluate our method through comprehensive quantitative and qualitative analyses. As shown in Table. I, our approach achieves state-of-the-art performance across three scenarios: MatrixCity, Rubble and Building. In the Building scene, our approach achieves substantial gains over Scaffold-GS, with margins of +0.048 (SSIM), +1.48 (PSNR), and +0.058 (LPIPS), underscoring its superior rendering quality. We attribute this superiority to our collaborative framework’s enhanced adaptability to spatially structured scenes - compared to Scaffold-GS’s hash grid encoding, our architecture models geometric priors more effectively. Notably, Pixel-GS exhibits marked performance degradation in this scenario, suggesting that pixel-optimized point cloud processing struggles with large-scale environments containing complex spatial structures. Fig. 3 presents visual comparisons between our method and existing 3DGS baselines. The experimental results show that our approach consistently preserves high-fidelity detail reconstruction while effectively mitigating artifacts common in other methods. Particularly in the Building scene, our method achieves more accurate reconstruction of fine structural details such as solar panel grid patterns, demonstrating clear advantages over competing approaches.
@@ V-D2 Result on Tanks & Temples dataset
We present the quantitative results on Tanks and Temples in Table. II. Our method demonstrates consistent superiority over other competitive approaches across all metrics on the ’train’ scene. Specifically, compared to Scaffold-GS, our approach achieves significant improvements of 0.024 in SSIM, 1.44 in PSNR, and 0.055 in LPIPS. As visually confirmed in Fig. 4, our method effectively separates the background sky from distant mountains in the rendered imagery. For the ’truck’ scene, our method exhibits superior performance in both PSNR and LPIPS metrics compared to other approaches. While it shows a marginal 0.005 deficit in SSIM relative to Scaffold-GS, our method produces visually superior results. Specifically, it better reconstructs fire hydrant contours and significantly reduces artifacts in the truck’s window regions.
@@ V-E2 Effectiveness of Cross-View Assisted Training Strategy
In this section, we conduct ablation studies on the proposed Cross-View Stereo Regularization (CVSR) module, Cross-View Augmented Densification (CVAD) module, and Cross-View Consistency Loss (CVCL) module. All modules are implemented using our cross-view assisted training strategy. Furthermore, we specifically optimize the CVCL module through pruning operations. As shown in Table. VI, integrating these key components enables the model to achieve optimal performance, with particularly significant improvements in SSIM metrics. Qualitative results in Fig. 8 demonstrate that incorporating these components significantly enhances rendering quality by improving both detail clarity (e.g., street lamp) and artifact suppression.

### v2

@@ I Introduction
In this work, we propose Structure–View Collaborative Gaussian Splatting (SplatCo), a unified framework for high-fidelity reconstruction and rendering of large-scale outdoor scenes. Unlike prior approaches that separately address geometry representation or view consistency, SplatCo integrates both global structure modeling and multi-view coherence into a single, jointly optimized system. We incorporate multi-scale contextual information into a global tri-plane backbone using a three-level compensation mechanism, which aligns coarse scene layout with detailed local geometry and texture. We also introduce a cross-view assisted training strategy that combines the gradient synchronization mechanism proposed in MVGS [13] with a structural-consistency-guided pruning mechanism, which simulatenously governs Gaussian densification and removes overfitted or inconsistent Gaussians based on structural consistency, to ensure consistent updates across overlapping views and stabilize Gaussian positioning and appearance during optimization. By combining these components in a unified training and inference pipeline, SplatCo achieves robust generalization and fine-detail reconstruction across complex, large-scale environments, as shown in Fig. 1.
@@ II-B Novel View Synthesis
3DGS has been widely adopted for modeling objects and scenes due to its speed and rendering quality. Readers may refer to recent surveys for comprehensive overviews of 3DGS and its applications [48, 49, 50]. While early applications focused on compact scenes, such as individual objects [51, 52], humans [53, 54, 55], and indoor environments [56, 57, 58], adapting 3DGS to large-scale outdoor scenes remains a challenge.
To bridge this gap, recent works propose scalable variants of 3DGS. Scaffold-GS [59] introduces a hierarchical, region-aware 3D representation via anchor points. DOGS [60] significantly improves training efficiency through a distributed paradigm that combines spatial partitioning with the ADMM solver. VastGaussian [61] and Hierarchy-GS [62] explore alternative partitioning strategies, each offering distinct architectural advantages. More recent developments include Octree-GS [12], which organizes Gaussian primitives in an octree structure to support adaptive multi-scale rendering, and CityGaussian [63], which incorporates level-of-detail techniques to optimize both training and rendering in urban-scale environments. Inspired by the explicit multi-view training strategy in MVGS [13], we combine it with the multi-level structure–view collaboration mechanism. Our approach requires both structural consistency constraints and cross-view consistency constraints. Computational results demonstrate that this combination is effective for achieving coherent rendering in large-scale unbounded scenes.
Despite these advances, two core limitations remain. First, rigid partitioning of space often breaks continuity of semantic structures, such as roads, facades or vegetation. Second, the lack of structural coordination and view-collaboration mechanisms limits the rendering fidelity required in large-scale scenarios. These challenges often lead to visual artifacts, such as ghosting, texture flickering, and inconsistent shading.
Figure 2: Overview. We propose a Cross-Structure Collaboration Module (CSCM) that hierarchically integrates tri-planes with context grids, effectively preserving fine-grained details while maintaining consistent global geometry. Additionally, we introduce a Cross-View Assisted Training Strategy (CVATS) that seamlessly integrates the Multi-View Augmented Densification (MVAD) and Multi-View Regulated Learning (MVRL) from MVGS [13] with the proposed Cross-View Consistent Loss (CVCL). This synergy enforces across overlapping cameras, effectively mitigating single-view overfitting in large-scale scenes.
@@ III-C Cross-Structure Collaboration Module
Tri-plane Feature Acquisition. Tri-Plane is a static 3D instance of the $k$-planes factorization [65], which represents the 3D space using three $m$-channel feature plane maps $\mathbf{F}\in\mathbb{R}^{w\times h\times m}$, each corresponding to one of the axis-aligned planes: $xy$, xzxz, and $yz$, with spatial resolution $w\times h$. To obtain fine-grained tri-plane features for a queried 3D point $\mathbf{p}$, we first apply a tri-plane attention module to generate attention-enhanced tri-plane features $\mathbf{F}^{TA}$. This module captures relative features across all three planes from the global tri-plane context, thereby enhancing the global perception capability compared to traditional tri-plane representations that focus solely on individual planes. The formulation is as follows:
@@
where $i$ denotes the axis-aligned planes: $xy$, xzxz, and $yz$, and $\text{TA}(\cdot)$ represents the tri-plane attention operation. The operator $||_{j\in\{xy,xz,yz\}}(\cdot)$ indicates concatenating feature planes along the channel dimension in the order of $xy$, xzxz, and $yz$. $\text{SA}(\cdot)$ and $\text{CA}(\cdot)$ denote the spatial attention and channel attention operations [66] applied to the feature maps, respectively. $\rho_{i}(\cdot)$ corresponds to the operation of extracting the tri-plane attention feature map for the dimension indexed by $i$ from the previously concatenated feature maps.
@@
where $G$ is the resolution grid that the query point $\mathbf{p}$ falls on, and $v$ represents the vertices of $G$. The notation $\{\cdot\}$ is used to compute the normalized absolute distance between the coordinates of two points along each axis, while $\cdot$] denotes the concatenation operation. $f_{p}$ refers to the inherent features of the query point, $O$ is the offset of the neural Gaussian center associated with the query point, and $s$ denotes the scales of these Gaussians.
@@ III-D Cross-View Assisted Training Strategy
To address this issue, we integrate the explicit multi-view regularization strategy proposed in MVGS [13] into our cross-view assisted training strategy. Specifically, we first employ a Multi-View Regulated Learning method from MVGS  [13] that leverages multiple views for collaborative optimization. Additionally, we also adopt the Multi-View Augmented Densification in MVGS [13] to improve point distribution by incorporating information from spatially distant views, which is particularly advantageous in large-scale scenes. Finally, a Cross-View Consistent Loss is introduced to enforce coherence between adjacent viewpoints and mitigate overfitting of Gaussians.
Multi-View Regulated Learning [13]. We extend the original single-view iterative training to a multi-view stereo approach that optimizes 3D Gaussians under multi-view supervision. In each iteration, we sample $M$ pairs of images and camera parameters, ensuring they are distinct from one another. We adopt the same multi-view regularization as in MVGS [13]:
@@
where $\mathcal{L}$ represents the loss at the current iteration, $G=\{G_{1},G_{2},\dots,G_{M}\}$ denotes the set of 3D Gaussians, with each $G_{M}$ being affected by large gradients from each view during multi-view training, and $E_{M}$ and $K_{M}$ are the extrinsic and intrinsic parameters of the camera, respectively.
@@
Multi-View Augmented Densification [13]. When optimizing across multiple viewpoints, particularly in large-scale scenes, selecting views that are too far apart is common. This results in limited overlap between perspectives, which increases the risk of overfitting Gaussian ellipsoids. To address this, we use the same distance-aware multi-view densification in MVGS [13]. This process involves normalizing the camera translations of sampled views to fit within a unit sphere, making our method adaptable to various scenes. Then it computes the relative translation distances $\{r_{i}|i=1,2,\dots,M\}$ between the cameras. The adaptive densification threshold with a self-adaptive threshold $\hat{\beta}$ in MVGS [13] is as follows:
@@
Mega-NeRF [CVPR22] [67]
0.553 24.06 0.516 0.547 20.93 0.504
Switch-NeRF [ICLR22] [68]
0.562 24.31 0.496 0.579 21.54 0.474
@@
III-E Update Gaussian Attributes &\& Rasterization
@@
where $\mathcal{A}_{i}$ represents the parameters of the $k$ Gaussian primitives for each anchor, and $\cdot$] denotes the concatenation operation.
@@ IV-C Fine-Grained Gaussian Regularization Loss
where $N_{g}$ represents the number of neural Gaussians in the scene, and $\prod(\cdot)$ indicates the product of the elements in a vector. In this case, it applies to the scale values $s_{i}$ of each Gaussian.
@@ V-A Datasets
[…]
@@ V-D1 Result on Mill-19 and MatrixCity dataset
We rigorously evaluate our method through comprehensive quantitative and qualitative analyses. As shown in Table. I, our approach achieves state-of-the-art performance across three scenarios: MatrixCity, Rubble and Building. In the Building scene, our approach achieves substantial gains over Scaffold-GS, with margins of +0.048 (SSIM), +1.48 (PSNR), and +0.058 (LPIPS), underscoring its superior rendering quality. We attribute this superiority to our collaborative framework’s enhanced adaptability to spatially structured scenes - compared to Scaffold-GS’s hash grid encoding, our architecture models geometric priors more effectively. Notably, Pixel-GS exhibits marked performance degradation in this scenario, suggesting that pixel-optimized point cloud processing struggles with large-scale environments containing complex spatial structures. For a fair comparison, we utilize the official open-source code of MVGS, which is based on the 3DGS framework. The sampled camera-image pair parameter $M$ in MVGS is also set to 4, maintaining consistency with our method. Our method demonstrably outperforms MVGS across the MatrixCity, Rubble, and Building scenes in most metrics. In the Rubble scene only, our method achieves an LPIPS score slightly below 0.012. The superior performance is primarily attributed to the Structure-View Collaboration Optimization approach, which mitigates the detail learning degradation issue in large-scale scenes,effectively preserving the ability to learn both the overall layout and local details. Fig. 3 presents visual comparisons between our method and existing 3DGS baselines. The experimental results show that our approach consistently preserves high-fidelity detail reconstruction while effectively mitigating artifacts common in other methods. Particularly in the Building scene, our method achieves more accurate reconstruction of fine structural details such as solar panel grid patterns, demonstrating clear advantages over competing approaches.
@@ V-D2 Result on Tanks & Temples dataset
We present the quantitative results on Tanks and Temples in Table. II. Our method demonstrates superiority over other competitive approaches across most metrics on the ’train’ scene. Specifically, compared to Scaffold-GS, our approach achieves significant improvements of 0.024 in SSIM, 1.44 in PSNR, and 0.055 in LPIPS. For a fair comparison, the camera-image pair sampling parameter $M$ in MVGS is also set to 4. Our method achieves performance gains over MVGS, yielding improvements of $0.007$ and $1.33$ on the SSIM and PSNR metrics, respectively. We also note that our method falls marginally behind MVGS on the LPIPS metric for the ’train’ and ’Truck’ scenes. As visually confirmed in Fig. 4, our method effectively separates the background sky from distant mountains in the rendered imagery. For the ’truck’ scene, our method exhibits superior performance in both PSNR and LPIPS metrics compared to other approaches. While it shows a marginal 0.005 deficit in SSIM relative to Scaffold-GS, our method produces visually superior results. Specifically, it better reconstructs fire hydrant contours and significantly reduces artifacts in the truck’s window regions.
@@ V-E2 Effectiveness of Cross-View Assisted Training Strategy
In this section, we conduct ablation studies on the MVRL module, MVAD module, and CVCL module. All modules are implemented using our cross-view assisted training strategy. Furthermore, we specifically optimize the CVCL module through pruning operations. As shown in Table. VI, integrating these key components enables the model to achieve optimal performance, with particularly significant improvements in SSIM metrics. Qualitative results in Fig. 8 demonstrate that incorporating these components significantly enhances rendering quality by improving both detail clarity (e.g., street lamp) and artifact suppression.